OpenStudy (anonymous):

what is the greatest common factors of 1?

OpenStudy (anonymous):

PLEASE HELP!!

OpenStudy (anonymous):

1 i think wait ill check it out

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yea it one because 1x itself = 1 and 1 times other number does not equal one

OpenStudy (anonymous):

for example The factors of 12 are 1,2,3,4,6 and 12 ... ... because 2 × 6 = 12, or 4 × 3 = 12, or 1 × 12 = 12.

OpenStudy (anonymous):

7, 20 • 3, 45 • 15, 18 • 4, 16 is it one of these?

OpenStudy (anonymous):

ummm hold on i think i have to check my math

OpenStudy (anonymous):

Multiply those factors both numbers have in common. If there are no common prime factors, the GCF is 1. there for it none of the above \

OpenStudy (anonymous):

i think it is 15/18

OpenStudy (anonymous):

ummm try it i dont went to give you the wrong answer

OpenStudy (anonymous):

no it's 7/20

OpenStudy (anonymous):

i think any way

OpenStudy (anonymous):

this is what i just got The Greatest Common Factor (GCF) of the numbers 1 is 1 because 1 is the greatest number that divides evenly into all of them.